# iVideoGPT: Interactive VideoGPTs are Scalable World Models

### v2

@@ 5 Related Work
Learning general world models in visual domains remains a significant challenge in model-based reinforcement learning. A straightforward method involves learning action-conditioned video prediction models [66, 43]. Advanced model-based RL algorithms [28, 30, 31, 77, 32] utilize latent imagination for more efficient and accurate rollouts but complicate themselves by tightly coupling model and policy learning. We show that this complexity can be reduced with powerful world models that have accumulated generalizable knowledge beyond specific tasks. Recent efforts to facilitate this include leveraging scalable architectures like transformers [60] and pre-training from large-scale data [95, 59]. Of particular relevance to our work are UniSim [98] and Genie [11], which have developed extensively trained world models with diffusion and masked models, respectively, though neither is publicly available. Our work distinguishes itself by utilizing a generic autoregressive transformer framework, advancing the flexibility of scalable world models.
@@
Recent developments in Internet-scale video generation models now enable the synthesis of realistic videos conditioned on class labels, text descriptions, and initial frames—the latter also known as the video prediction problem. Various models have been developed, including deterministic RNNs [80, 92], variational autoencoders [17, 2, 29, 3], diffusion [36, 10], masked [100, 26], and autoregressive models [97, 48, 53]. However, most recent works do not treat video prediction as a dynamics modeling problem and perform spatiotemporal compression [97, 7], thus providing limited interactivity to serve as world models. We achieve both compressive tokenization and interactivity by context-aware representation, employing cross-attention mechanisms with minimal inductive bias. This method diverges from previous techniques that rely on motion vectors [41] or optical flows [50] and offers a more generic form of video tokenization.
@@ 6 Discussion
While iVideoGPT marks significant progress, there is substantial room for improvement. We found limited diversity in publicly available robotic data, including the large-scale Open X-Embodiment dataset, and initiated efforts to transfer knowledge from human videos [24]. We believe iVideoGPT should be pre-trained on more extensive data [25] to bridge knowledge between humans and robots. This requires iVideoGPT to incorporate more modalities, such as multi-view observations, proprioceptive robot states, and actions, within the unified formulation beyond action-free video prediction. Specifically, to process high-dimensional visual observations, our compressive tokenization assumes that initial frames provide sufficient contexts for future frames, which works for low-level control tasks as model-based agents often foresee tens of steps, but may falter in scenarios with long videos and significant camera motion. This issue can be mitigated by keyframe extraction [49] but leaves an important future avenue of exploration.
@@ A.1 Architecture
Learning rate $5\times 10^{-4}$ $5\times 10^{-4}$ $5\times 10^{-4}$ $5\times 10^{-4}$ $5\times 10^{-4}$ $1\times 10^{-4}$
@@ A.2 Action-free Video Pre-training
We pre-trained iVideoGPT using 35 datasets from the Open X-Embodiment Dataset (OXE) [67] and Something-Something-v2 (SSv2) [24]. To construct our training dataset from OXE, we implement a filtering and weighting process similar to Octo [64]. Initially, we exclude datasets lacking image streams and those derived from mobile robots. Subsequently, datasets exhibiting excessive repetition or possessing low image resolutions are eliminated. The remaining datasets were categorized as either "large" or "small," and each was assigned a weight based on its size and diversity. For a comprehensive breakdown of the mixture, refer to Table 4. We selected $1\%$ of samples from each subset as validation data and used the rest for training.
@@
Something-Something-V2 [24] 193,690 1 15.0%
Total 1,491,063 - 100.0%
@@ A.4 Visual Planning
A.5 Visual Model-based Reinforcement Learning
Environments.
Meta-world [101], following MIT License, is a benchmark of 50 robotic manipulation tasks. We select six tasks for our experiments: Button Press Topdown Wall, Plate Slide, Hammer, Door Lock, Handle Pull Side, and Coffee Push. We set the maximum episode length to 200 environment steps with an action repeat of 2 across these tasks and adjust the number of training steps to match the varying difficulty levels. During experiments, we observed that high rewards do not consistently correlate with high success rates in the original Meta-world implementation. This discrepancy presents a challenge to the learning stability of agents. To address this, we introduce an additional bonus for task success $r_{\text{bonus}}=10.0$ alongside the original task reward $r_{\text{task}}$:
Implementation Details.
Table 5: Hyperparameters of model-based RL with iVideoGPT.
Model-based RL Hyperparameter Value
Model rollout Init rollout batch size 640
Interval 200 env. steps
Batch size 32
Horizon 10
Model training Init training steps 1000
Tokenizer training interval 40 env. steps
Transformer training interval 10 env. steps
Batch size 16
Sequence length 12
Context frames 2
Sampled future frames (tokenizer) 5
Learning rate $1\times 10^{-4}$
Weight decay 0
Optimizer Adam
Model-based RL Real data ratio 0.5
@@ A.5 Visual Model-based Reinforcement Learning
Appendix B Additional Experimental Results
@@ B.1 Qualitative Evaluation
We present additional examples of video predictions by iVideoGPT on various datasets in Figures 10, 12, 13, 14, 15, and 16. We also include an additional showcase of zero-shot predictions by the pre-trained transformer in iVideoGPT in Figure 11, supplementing Figure 7 of the main text.
Figure 10: Additional qualitative evaluation on the Open X-Embodiment dataset.
@@
B.2 Failure Case Analysis for Visual Planning
@@
Our model performs sub-optimally on the RoboDesk open slide task from the VP2 benchmark, and in this section, we investigate the underlying causes.
@@
Figure 17: Failure case analysis on the RoboDesk open slide task from the VP2 benchmark, where, due to the low resolution of observations, our model fails to discriminate between subtle changes, particularly whether the robot’s gripper has made contact with the slide handle.
Appendix C Computational Resources

### v3

@@ 5 Related Work
Learning general world models in visual domains remains a significant challenge in model-based reinforcement learning. A straightforward method involves learning action-conditioned video prediction models [69, 45]. Advanced model-based RL algorithms [29, 31, 32, 80, 34, 33] utilize latent imagination for more efficient and accurate rollouts but complicate themselves by tightly coupling model and policy learning. We show that this complexity can be reduced with powerful world models that have accumulated generalizable knowledge beyond specific tasks. Recent efforts to facilitate this include leveraging scalable architectures like transformers [63] and pre-training from large-scale data [99, 62]. Of particular relevance to our work are UniSim [102] and Genie [12], which have developed extensively trained world models with diffusion and masked models, respectively, though neither is publicly available. Our work distinguishes itself by utilizing a generic autoregressive transformer framework, advancing the flexibility of scalable world models.
@@
Recent developments in Internet-scale video generation models now enable the synthesis of realistic videos conditioned on class labels, text descriptions, and initial frames—the last one also known as the video prediction problem. Various models have been developed, including deterministic RNNs [84, 96], variational autoencoders [18, 3, 30, 4], diffusion [38, 11], masked [104, 27], and autoregressive models [101, 50, 55]. However, most recent works do not treat video prediction as a dynamics modeling problem and perform spatiotemporal compression [101, 8], thus providing limited interactivity to serve as world models. We achieve both compressive tokenization and interactivity by context-aware representation, employing cross-attention mechanisms with minimal inductive bias. This method diverges from previous techniques that rely on motion vectors [43] or optical flows [52] and offers a more generic form of video tokenization.
@@ 6 Discussion
While iVideoGPT marks significant progress, there is substantial room for improvement. We found limited diversity in publicly available robotic data, including the large-scale Open X-Embodiment dataset, and initiated efforts to transfer knowledge from human videos [25]. We believe iVideoGPT should be pre-trained on more extensive data [26] to bridge knowledge between humans and robots. This also requires iVideoGPT to incorporate more modalities, such as multi-view observations, proprioceptive robot states, and actions, within the unified formulation beyond action-free video prediction. Specifically, to process high-dimensional visual observations, our compressive tokenization assumes that initial frames provide sufficient contexts for future frames, which works for low-level control tasks as model-based agents often foresee tens of steps, but may falter in scenarios with long videos and significant camera motion. This issue can be mitigated by keyframe extraction [51] but leaves an important future avenue of exploration. Finally, extending to more complex real-robot tasks is essential, as the benefits of model scaling to even larger sizes remain unobserved in this work within visually simple simulation for downstream control tasks.
Acknowledgements
We would like to thank many colleagues, in particular Yuchen Zhang, Lanxiang Xing, and Haixu Wu, for their valuable discussion. This work was supported by the National Natural Science Foundation of China (U2342217 and 62022050), the BNRist Project, the Huawei Innovation Fund, and the National Engineering Research Center for Big Data Software.
@@ A.1 Architecture
Learning rate $5\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $5\times 10^{-4}$ $1\times 10^{-4}$
@@ A.2 Action-free Video Pre-training
We pre-train iVideoGPT using 35 datasets from the Open X-Embodiment Dataset (OXE) [70] and Something-Something-v2 (SSv2) [25]. To construct our training dataset from OXE, we implement a filtering and weighting process similar to Octo [67]. Initially, we exclude datasets lacking image streams and those derived from mobile robots. Subsequently, datasets exhibiting excessive repetition or possessing low image resolutions are eliminated. The remaining datasets were categorized as either "large" or "small," and each was assigned a weight based on its size and diversity. We select $1\%$ of samples from each subset as validation data and use the rest for training. For SSv2, we manually select 95 classes with clear motion trends from the original 174 video classes as our pre-training data with a weighting of 15%. We use the official splits of SSv2 for training and validation. For a comprehensive breakdown of the mixture, refer to Table 4.
@@
Goal-conditioned prediction.
To train a goal-conditioned variant of iVideoGPT on the same dataset, we first fine-tune the previously obtained tokenizer using two randomly sampled frames as context for 550k training steps. Then, we train a transformer from scratch with the rearranged frame segment $\tilde{o}_{1:T}=(o_{T},o_{1},o_{2},\dots,o_{T-1})$ for 1 million steps. The architecture and training procedures remain consistent with the above setup.
@@
Something-Something-V2 [25] 120,581 1 15.0%
Total 1,417,954 - 100.0%
@@ A.4 Visual Planning
A.5 Visual Model-based RL
@@
Environments.
Meta-world [105], following MIT License, is a benchmark of 50 robotic manipulation tasks. We select six tasks for our experiments: Button Press Topdown Wall, Plate Slide, Hammer, Door Lock, Handle Pull Side, and Coffee Push. We set the maximum episode length to 200 environment steps with an action repeat of 2 and a frame stack of 3 across these tasks and adjust the number of training steps to match the varying difficulty levels. During experiments, we observed that high rewards do not consistently correlate with high success rates in the original Meta-world implementation. This discrepancy presents a challenge to the learning stability of agents. To address this, we introduce an additional bonus for task success $r_{\text{bonus}}=10.0$ alongside the original task reward $r_{\text{task}}$:
Implementation Details.
Baselines.
To compare our method with DreamerV3, which lacks native pre-training support, we use APV [81]—a method enabling action-free pre-training on DreamerV2—as a baseline, modified to incorporate DreamerV3 features. We pre-train this model on the same dataset as iVideoGPT.
Table 5: Hyperparameters of model-based RL with iVideoGPT.
Model-based RL Hyperparameter Value
Model rollout Init rollout batch size 640
Interval 200 env. steps
Batch size 32
Horizon 10
Model training Init training steps 1000
Tokenizer training interval 40 env. steps
Transformer training interval 10 env. steps
Batch size 16
Sequence length 12
Context frames 2
Sampled future frames (tokenizer) 5
Learning rate $1\times 10^{-4}$
Weight decay 0
Optimizer Adam
Model-based RL Real data ratio 0.5
Appendix B Extended Experimental Results
@@ B.1 Qualitative Evaluation
We present additional examples of video predictions by iVideoGPT on various datasets in Figures 12, 13, 15, 16, 17, 18, and 19. We also include an additional showcase of zero-shot predictions by the pre-trained transformer in iVideoGPT in Figure 14, supplementing Figure 8 of the main text.
Additionally, we showcase prediction examples from the large-scale pre-trained DreamerV3-XL on the Open X-Embodiment dataset in Figures 11.
Figure 11: Qualitative evaluation on the Open X-Embodiment dataset for Dreamerv3-XL pre-trained on the same dataset as ours.
Figure 12: Additional qualitative evaluation on the Open X-Embodiment dataset for action-free video prediction.
Figure 13: Additional qualitative evaluation on the Open X-Embodiment dataset for goal-conditioned video prediction.
@@
B.2 Human Study
Numerical metrics like FVD don’t always align with human-judged visual quality. To address this, we conduct a human user study on the prediction results of various models. Due to the lack of official pretrained models for most baselines, we are only able to compare iVideoGPT with VideoGPT [101] and MCVD [93] on the action-free BAIR dataset. We generate videos using these three models from the test set and ask users to label preferences between two randomly sampled videos based on the physical naturalness and feasibility of robot-object interactions. A total of 386 annotations are collected from 9 participants. The results in Figure 20 demonstrate that iVideoGPT is preferred by human annotators more.
Figure 20: Human study. Videos generated by three models, VideoGPT, MCVD, and iVideoGPT, on the action-free BAIR dataset are presented to human users, who label their preferences based on the physical naturalness and feasibility of robot-object interactions.
B.3 Visual Planning
Quantitative results on the VP2 benchmark are reported in Table 6.
Table 6: Quantitative results on the VP2 benchmark, reporting mean, min, and max performance over various control runs.
Tasks
Success
rate
iVideoGPT
[…]
Struct-
VRNN
Simulator
Robosuite push mean 0.7833 0.6760 0.7980 0.7733 0.8260 0.5540 0.9350
max 0.7950 0.7900 0.8400 0.7900 0.8500 0.6000 0.9500
min 0.7750 0.6400 0.7600 0.7400 0.7900 0.5000 0.9200
Flat block mean 0.0333 0.0917 0.0200 0.0500 0.0400 0.0467 0.1333
max 0.0417 0.1333 0.0333 0.0667 0.1000 0.1333 0.1333
min 0.0250 0.0667 0.0000 0.0333 0.0000 0.0000 0.1333
Open drawer mean 0.3750 0.2533 0.1667 0.1167 0.0400 0.0267 0.7667
max 0.3917 0.3333 0.2667 0.1333 0.1000 0.1000 0.7667
min 0.3500 0.1333 0.0667 0.1000 0.0000 0.0000 0.7667
Open slide mean 0.1611 0.3533 0.5733 0.1833 0.0867 0.1267 0.7167
max 0.1917 0.4000 0.7333 0.2000 0.1667 0.2333 0.7333
min 0.1250 0.2667 0.4667 0.1667 0.0333 0.0667 0.7000
Blue button mean 0.9556 0.9400 0.9733 0.9500 0.9467 0.8667 1.0000
max 0.9833 0.9667 1.0000 1.0000 0.9667 0.9000 1.0000
min 0.9333 0.8667 0.9333 0.9000 0.9333 0.8000 1.0000
Green button mean 0.8250 0.8400 0.8133 0.8333 0.6400 0.6800 0.9667
max 0.8667 0.9000 0.9000 0.8333 0.7000 0.8000 0.9667
min 0.7833 0.7667 0.7667 0.8333 0.6000 0.5667 0.9667
Red button mean 0.9222 0.5867 0.7600 0.7333 0.2400 0.3067 0.9000
max 0.9333 0.6333 0.8667 0.7333 0.3333 0.3333 0.9000
min 0.9000 0.5000 0.6333 0.7333 0.1333 0.2333 0.9000
Upright block mean 0.4472 0.5133 0.4867 0.5667 0.6200 0.3333 0.9000
max 0.4667 0.5667 0.6667 0.6000 0.7333 0.3667 0.9000
min 0.4250 0.5000 0.4000 0.5333 0.5000 0.3000 0.9000
B.4 Visual Model-based RL
Comparison to FitVid-based world models.
Although FitVid [4] is originally designed for the video prediction task and has not been used as world models for MBRL, we have implemented a baseline using FitVid by replacing iVideoGPT in our implementation. To predict rewards, we add an MLP head on top of FitVid’s latent states, parallel to the image decoder. As shown in Figure 21, MBPO with iVideoGPT outperforms FitVid on 5 out of 6 tasks and performs comparably on the remaining one. We also qualitatively observe that FitVid’s imagined trajectories are blurrier compared to ours, which hinders its ability to simulate real environments accurately and may hurt MBRL performance.
B.5 Computational Efficiency
We report training and inference time and memory usage with various tokenizers in Table 7 and 8, respectively. Our proposed compressive tokenization provides significant memory savings during training and faster rollouts during generation. We note that although we use a more complicated tokenizer design, it is not the bottleneck of generation time. Additionally, although we use more tokens for context frames compared to the $4\times 4$ tokenizer, generation time is primarily influenced by the number of forward passes of the autoregressive transformer, which remains the same.
Table 7: Training efficiency of the transformer with various tokenizers, measured on 40G A100 GPUs with a per-device batch size of 16.
Tokenizer Speed (#iters/sec) Memory (GB)
$4\times 4$ 3.10 10.6
$16\times 16$ N/A OOM
Ours 2.62 22.3
Table 8: Generation efficiency with various tokenizers, measured on an RTX 4090 GPU with a batch size of 1.
Tokenizer Tokenize (sec) Generation (sec) Detokenize (sec) Memory (GB)
$4\times 4$ 0.27 1.13 0.05 1.98
$16\times 16$ 0.26 22.5 0.04 1.90
Ours 0.29 1.11 0.06 2.33
Figure 21: Visual model-based RL on Meta-world, comparing to an additional baseline using FitVid as world models.
Appendix C Extended Discussions
C.1 Differences with IRIS
Discrete tokenization and autoregressive transformers are prevalent in contemporary deep learning due to their simplicity and generality. iVideoGPT generally shares this architecture with IRIS [63], but possesses distinguishing features, summarized as follows:
  • Pre-training and fine-tuning paradigm: iVideoGPT is designed for a paradigm that involves pre-training on large-scale videos and fine-tuning on various downstream tasks. In contrast, IRIS focuses solely on MBRL with Transformer-based world models trained from scratch in the Atari domain.
  • Efficient tokenization: iVideoGPT proposes novel compressive tokenization to significantly reduce the number of tokens, saving time and memory (see Table 7 and 8), while IRIS uses per-frame tokenization.
  • Flexible action-conditioning design: iVideoGPT employs slot tokens with optional additive action embeddings to support both action-free pre-training and action-conditioned fine-tuning, while IRIS strictly treats discrete Atari actions as tokens.
  • Off-policy MBRL implementation: iVideoGPT uses an off-policy RL algorithm while IRIS performs on-policy learning. On-policy learning needs a large number of model rollouts, which, combined with inefficient tokenization, results in  7 days for 100k-environment-step training. In comparison, iVideoGPT only needs $\sim$4 hours.
C.2 Differences with VideoGPT
We elaborate on the the difference between the tokenizer in VideoGPT [101] and ours, and how they impact interactivity.
VideoGPT uses a VQVAE for video that relies on a series of 3D convolutions to downsample across space and time. For example, it downsamples original pixels from $16\times 64\times 64$ to discrete tokens of $8\times 32\times 32$ or $4\times 16\times 16$, depending on the downsampling ratio. The key issue is that this non-causal downsampling over the temporal dimension results in each token containing information from a window of frames. As a result, the entire video of a fixed length can only be reconstructed after VideoGPT generates all tokens. As shown in Figure 2, VideoGPT only allows the input of future action sequences at the beginning of prediction, preventing an agent from interactively determining its actions based on predicted observations. In contrast, our tokenizer discretizes video frames separately, using a conditional mechanism to handle temporal redundancy, enabling frame-by-frame video generation and allowing for intermediate action intervention.
Moreover, our tokenizer’s novel design, with its cross-attention mechanism, is more efficient in handling temporal redundancy, converting videos into significantly fewer tokens ($L=511$ with $N=256,n=16,T=16,T_{0}=1$ as stated in Section 3.1). In contrast, VideoGPT finds that using a larger downsampling ratio than a token size $8\times 32\times 32$, results in worse performance.
C.3 Failure Case Analysis for Visual Planning
@@
Our model performs sub-optimally on the RoboDesk open slide task from the VP2 benchmark. In this section, we investigate the underlying causes through case studies, attributing the performance issues to limitations in both our model and the benchmark.
Inaccurate model prediction.
@@
Imperfect built-in reward design.
We observe that no current model in the VP2 benchmark consistently outperforms other models across all tasks, and iVideoGPT is no exception. Beyond models’ inaccuracies in prediction for severely out-of-distribution (OOD) actions, our analysis of this inconsistent performance also reveals flaws in the benchmark’s built-in reward design.
In VP2, scores for sampled actions are estimated mainly by a learned classifier that assesses task success based on model-predicted frames. This classifier, trained by the VP2 authors, appears to lack robustness and is easily fooled by OOD inputs, assigning high rewards to low-quality or unlikely-to-succeed predicted trajectories (see examples in Figure 23). Such an imperfect reward function likely contributes to the mixed results observed on this benchmark, with iVideoGPT even outperforming the oracle simulator in one task. Addressing visual planning with imperfect rewards is an independent research problem and beyond the scope of this paper.
Figure 22: Failure case analysis on the RoboDesk open slide task from the VP2 benchmark, where, likely due to the low resolution of observations, our model fails to discriminate between subtle changes, particularly whether the robot’s gripper has made contact with the slide handle.
Figure 23: Imperfect built-in reward in VP2 benchmark. A learned reward model can assign high rewards to predicted transitions that are less likely to succeed, which can mislead optimizers in model-predictive control.
Appendix D Computational Resources
@@
Appendix E Broader Impact
World models advance the development of autonomous machine intelligence, particularly through the valuable visual insights offered by videos. However, their full potential remains untapped without scalable and interactive architectures capable of distilling vast amounts of commonsense knowledge from multimodal data. This paper, we believe, takes an important step by introducing a flexible framework with a specific focus on the robotic manipulation domain. Our results may pave the way for higher-quality world models applicable across diverse domains, enhancing performance in control tasks of embodied intelligence. Despite the benefits, designing and training these models is challenging, requiring substantial computational power and increasing the carbon footprint. Using underdeveloped, inaccurate world models for autonomous control could lead to risky actions and potential physical damage, which can be mitigated by developing uncertainty-aware models to prevent uncertain actions. Conversely, the advancement of these models could lead to job displacement in sectors relying on manual control tasks. Additionally, the underlying techniques for world models can be misused to generate synthetic videos that mimic real events or people. However, since our model is merely a research prototype, trained only with robotic and human manipulation data and relatively small in scale, we do not anticipate immediate negative societal impacts such as deepfakes or job displacement.